OpenStudy (anonymous):

Jeffrey is training to run a marathon. On Monday he ran and walked a total of 11 miles. If he ran 2 miles more than twice as far as he walked, how many miles did he run?

OpenStudy (anonymous):

x - how much did he run y - how much did he walk x+y = 11 -> y = 11-x 2y = x+2 2*(11-x) = x+2 22-2x = x+2 -3x = -20 x=20/3 y = 11-x y = 11 - (20/3) y = (33/3) - (20/3) y = 13/3 he ran 20/3 miles and walked 13/3 miles